OpenStudy (anonymous):

ABC is a right triangle. If AC=4 and BC=10, find AB. leave your answer in simplest radical form. (please explain how to leave your answer in simplest radical form?)

Parth (parthkohli):

She didn't ask the names of every single person alive @katysmiles123 she just wants to know how to simplify a radical

Parth (parthkohli):

Factorize the number inside the radical, and then cancel the ones repeating two times, and bring that outside the radical.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{8} = \sqrt{2 \times 2 \times 2} = \sqrt{\cancel{2 \times 2} \times 2} = 2\sqrt2 }\) Example

OpenStudy (anonymous):

How would i get the first answer? i don't even know how to do that, then can you explain to me the whole radical thing?

OpenStudy (anonymous):

a picture would help, because you have a right triangle, but we don't know if 4 and 10 are the sides of the legs, or if 10 is the hypotenuse

OpenStudy (anonymous):

if 10 is the length of the hypotenuse, then \[4^2+y^2=10^2\] but if 10 is the length of the other leg, then \[4^2+10^2=y^2\]

Parth (parthkohli):

Pythagorean theorem. the missing side is either a leg or the hypotenuse.

Parth (parthkohli):

^ yes

OpenStudy (anonymous):

BC is the hypotenuse

OpenStudy (anonymous):

it will not let me post my drawing. i dunno why.